The thermite reaction is very exothermic, the heat of the reaction sufficient to melt the iron that is produced by the reaction:

2Al + Fe2O3 Al2O3 + 2Fe H=?

a) calculate the standard enthalpy change for this reaction?
b) if 40.0 g of iron (III) oxide we’re reacted with excess aluminum, how much heat would be produced ?
c) of the heat produced was used to heat 2.00L of water at 23.0C, what would the final temperature of the water be?
1
Expert's answer
2019-07-16T07:12:26-0400

2Al(s)+Fe2O3(s) -> Al2O3(s) + 2Fe(s)

a) ΔH = ΣΔH(products) -ΣΔH(reactants)

ΔH (Al) = 0

ΔH(Fe2O3) = -824.4kJ/mol

ΔH (Fe) = 0

ΔH (Al2O3) = -1676 kJ/mol

ΔH = (-1676+0*2)-(-824.2+0*2) =-851.8 kJ


b) n=m/M

n(Fe2O3) = 40.0 / 159.69 =0.250 mol

according to equation 1 mole of Fe2O3 gives 851.8 kJ of heat

we have 0.250 mole of Fe2O3 which give x kJ of heat

1/0.250 = 851.8/x

x = 213.4 kJ


c) q = cm(T2-T1)

213400 = 4184*2(T2 - 23)

T2 = 48.5 C